# Test-time regression: a unifying framework for designing sequence models with associative memory

### v1

@@ Appendix A Bounding the norm of a linear regression layer
Let $\mathbf{y}_{t}=\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\mathbf{q}_{t}$ and let $\lVert\mathbf{A}\rVert_{2}$ indicate the spectral norm of a matrix $\mathbf{A}$, a matrix norm induced by the L2 norm on vectors. Then
$\lVert\mathbf{y}_{t}\rVert_{2}$ $=\lVert\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\mathbf{q}_{t}\rVert_{2}$ (41)
$\leq\lVert\mathbf{q}_{t}\rVert_{2}\sup_{\lVert\mathbf{q}\rVert_{2}=1}\lVert\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\mathbf{q}\rVert_{2}$ (42)
$=\lVert\mathbf{q}_{t}\rVert_{2}\lVert\mathbf{V}{}^{\top}\mathbf{K}(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\rVert_{2}$ (43)
$\leq\lVert\mathbf{q}_{t}\rVert_{2}\lVert\mathbf{V}{}^{\top}\mathbf{K}\rVert_{2}\lVert(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\rVert_{2}$ (44)
$=\lVert\mathbf{q}_{t}\rVert_{2}\lVert\sum_{i=1}^{t}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}\rVert_{2}\lVert(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\rVert_{2}$ (45)
$\leq\lVert\mathbf{q}_{t}\rVert_{2}\left(\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}\right)\lVert(\mathbf{K}{}^{\top}\mathbf{K})^{-1}\rVert_{2}$ (46)
$=\lVert\mathbf{q}_{t}\rVert_{2}\left(\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}\right)\max_{i}\lambda_{i}((\mathbf{K}{}^{\top}\mathbf{K})^{-1})$ (47)
$=\lVert\mathbf{q}_{t}\rVert_{2}\left(\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}\right)\max_{i}\frac{1}{\lambda_{i}(\mathbf{K}{}^{\top}\mathbf{K})}$ (48)
$=\frac{\lVert\mathbf{q}_{t}\rVert_{2}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}}{\lambda_{\text{min}}(\mathbf{K}{}^{\top}\mathbf{K})}.$ (49)
Hence $\lVert\mathbf{y}_{t}\rVert_{2}\leq\frac{\lVert\mathbf{q}_{t}\rVert_{2}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}}{\lambda_{\text{min}}(\mathbf{K}{}^{\top}\mathbf{K})}$. When we approximate $\mathbf{K}{}^{\top}\mathbf{K}\approx\mathbf{I}$, as in linear attention, we lose the self-normalizing property of dividing by ${\lambda_{\text{min}}(\mathbf{K}{}^{\top}\mathbf{K})}$ since the denominator becomes 1. This explains how output normalization by Qin et al. (2022) is an attempt at restoring this intrinsic self-normalizing property of linear regression.

### v3

@@ Appendix A Bounding the norm of a linear regression layer
Let $\mathbf{y}_{t}=\mathbf{V}_{t}{}^{\top}\mathbf{K}_{t}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1}\mathbf{q}_{t}$ and let $\lVert\mathbf{A}\rVert_{2}$ indicate the spectral norm of a matrix $\mathbf{A}$, a matrix norm induced by the L2 norm on vectors. Then
$\lVert\mathbf{y}_{t}\rVert_{2}$ $=\lVert\mathbf{V}_{t}{}^{\top}\mathbf{K}_{t}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1}\mathbf{q}_{t}\rVert_{2}$ (59)
$\leq\lVert\mathbf{q}_{t}\rVert_{2}\sup_{\lVert\mathbf{q}\rVert_{2}=1}\lVert\mathbf{V}_{t}{}^{\top}\mathbf{K}_{t}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1}\mathbf{q}\rVert_{2}$ (60)
$=\lVert\mathbf{q}_{t}\rVert_{2}\lVert\mathbf{V}_{t}{}^{\top}\mathbf{K}_{t}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1}\rVert_{2}$ (61)
$\leq\lVert\mathbf{q}_{t}\rVert_{2}\lVert\mathbf{V}_{t}{}^{\top}\mathbf{K}_{t}\rVert_{2}\lVert(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1}\rVert_{2}$ (62)
$=\lVert\mathbf{q}_{t}\rVert_{2}\lVert\sum_{i=1}^{t}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}\rVert_{2}\lVert(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1}\rVert_{2}$ (63)
$\leq\lVert\mathbf{q}_{t}\rVert_{2}\left(\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}\right)\lVert(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1}\rVert_{2}$ (64)
$=\lVert\mathbf{q}_{t}\rVert_{2}\left(\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}\right)\max_{i}\lambda_{i}((\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})^{-1})$ (65)
$=\lVert\mathbf{q}_{t}\rVert_{2}\left(\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}\right)\max_{i}\frac{1}{\lambda_{i}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})}$ (66)
$=\frac{\lVert\mathbf{q}_{t}\rVert_{2}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}}{\lambda_{\text{min}}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})}.$ (67)
Hence $\lVert\mathbf{y}_{t}\rVert_{2}\leq\lVert\mathbf{q}_{t}\rVert_{2}\sum_{i=1}^{t}\lVert\mathbf{v}_{i}\rVert_{2}\lVert\mathbf{k}_{i}\rVert_{2}/\lambda_{\text{min}}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})$. When we approximate $\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t}\approx\mathbf{I}$, as in linear attention, we lose the self-normalizing property of dividing by ${\lambda_{\text{min}}(\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t})}$ since the denominator becomes 1. This explains how output normalization by Qin et al. (2022) is an attempt at restoring this intrinsic self-normalizing property of linear regression.